Ships A and B leave port together. For the next two hours, ship A travels at 22 mph in a direction 40 ∘west of north while ship B travels 28 ∘ east of north at 29 mph .

What is the distance between the two ships two hours after they depart?

What is the speed of ship A as seen by ship B?
Answer #1

rax = - 22 * 2 sin 40 = - 28.28

ray = 22 *2 *cos 40 = 33.7

rbx = 29* 2 sin 28 = 27.23

rby = 29 * 2 cos 28 = 51.21

a)

now distance between the ships

d^2 = ( 51.21 - 33.7)^2 + (27.23 + 28.28)^2

d = 58.206 mi

=====

b)

reltive velocity

v = 58.206 / 2 = 29.1 mph

=====

Comment before rate in case any doubt, will reply for sure.. goodluck